OpenStudy (anonymous):

Please help me with this: Approximately 9% of U.S. children have asthma. Consider an elementary school with 500 children. 1. What is the probability that AT LEAST 45 children have asthma? 2. What is the probability that AT MOST 39 children have asthma? 3. What is the probability that AT MOST 51 children have asthma?

OpenStudy (accessdenied):

Interesting problem -- it looks like you use binomial distribution, but I'm not 100% sure on that.

OpenStudy (anonymous):

Ok. Thanks for trying. Are you familiar with margin of error?

OpenStudy (anonymous):

Oh, Okay. Well thank you for trying. (:

OpenStudy (accessdenied):

Maybe @FoolForMath would have a better idea about this... D:

OpenStudy (anonymous):

Poisson's formula?

OpenStudy (anonymous):

It deals with P= probability & N=trials

jimthompson5910 (jim_thompson5910):

Use the binomial probability distribution formula P(X = x) = (n C x)*((p)^x)*(1-p)^(n-x) This formula works for small values of x when it comes to computing cdf values (since you're adding individual values to compute the cdf) However, larger x values make things tedious. So it's better to use a calculator. One can be found here: http://stattrek.com/online-calculator/binomial.aspx

OpenStudy (anonymous):

How would i plug in my numbers?

jimthompson5910 (jim_thompson5910):

Well since x = 45, you need to add up 45 different values to find the cdf...which is way too tedious. So it's better to use the calculator. Follow that link and type in 0.09 for the probability (since 9% have it), 500 for the number of trials, and 45 for the value of x When you hit "calculate", it'll give you every bit of information concerning 45 successes The piece of info you want is P(X >= 45) which will be the last box This is all of course for #1. You'd do the same (but with different numbers) for #2 and #3

OpenStudy (anonymous):

Ok, so it gives me like $ answers, which one do i use?

jimthompson5910 (jim_thompson5910):

for #1, you want the last answer because you want to find P(X >= 45) (since you want AT LEAST 45)

jimthompson5910 (jim_thompson5910):

for #2 and #3, you want the second box answer (use different numbers of course)

OpenStudy (anonymous):

Oh. Ok, thank you SOO much. (: So would 57% be 0.57 or 0.057?

jimthompson5910 (jim_thompson5910):

which problem are you referring to, #1, #2, or #3?

jimthompson5910 (jim_thompson5910):

57% is 0.57 in decimal form

OpenStudy (anonymous):

Thank you. Its on a complete different problem. just making sure of that.

jimthompson5910 (jim_thompson5910):

oh ok, just making sure

OpenStudy (anonymous):

So for problem #2, the seconf answer is 0.154 The only multiple choice answers i have are 0.16 0.34 0.84 1

jimthompson5910 (jim_thompson5910):

hmm that's interesting...maybe there's a typo somewhere or a roundoff error. The closest thing is 0.16, so that's my best guess.

OpenStudy (anonymous):

The Same thing for my Other problems.

jimthompson5910 (jim_thompson5910):

for #3 as well? what are your choices there?

OpenStudy (anonymous):

0.135 0.16 0.34 0.84

OpenStudy (anonymous):

I will skip that. Do you know anything about Margin of Error?

jimthompson5910 (jim_thompson5910):

it looks like the answer to #3 is 0.84

jimthompson5910 (jim_thompson5910):

yes, what do you want to know about it

OpenStudy (anonymous):

My paper says: Find the margin of error for a survey that has the given sample size. Round your answer to the nearest tenth of a percent: 1. 225 2. 100 3. 625 4. 3600 5. 4200 6. 776 7. 390 8. 8000 & my answer box has answers like: + or - 1.1% , + or - 3.5% How do i do this?

jimthompson5910 (jim_thompson5910):

Are you given a standard deviation?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

It says something about use this formula: + or - 1/squareroot of N

jimthompson5910 (jim_thompson5910):

so then the margin of error for #1 is ME = +- 1/sqrt(N) ME = +- 1/sqrt(225) ME = +- 1/15 ME = +- 0.0667 ME = +- 6.67% ME stands for the Margin of Error and +- is shorthand for + or - You'd do the same for the rest.

OpenStudy (anonymous):

Ok so, number 2 would be: 0.1 so, +-1.1% ?

jimthompson5910 (jim_thompson5910):

yes you nailed it

OpenStudy (anonymous):

(: Thank you. can you help me with the last one, 8000. i got the answer 0.0111803399

jimthompson5910 (jim_thompson5910):

You got that one as well, so the margin of error is +-1.1%

OpenStudy (anonymous):

Thank youuu (:

jimthompson5910 (jim_thompson5910):

oh wait, I'm not thinking, the margin of error for #2 is +-10%, not +-1.1%...sorry about that

OpenStudy (anonymous):

oh ok. so # 4 3600 would be: +-1.7%

jimthompson5910 (jim_thompson5910):

yes since 1/sqrt(3600) = 0.017 which converts to 1.7%

OpenStudy (anonymous):

& number 5 would be 0.015430335 so, +-1.5%

jimthompson5910 (jim_thompson5910):

yes it is, nice work

OpenStudy (anonymous):

Ok, im stuck on this one. 1/SR390 = 0.0506369684

jimthompson5910 (jim_thompson5910):

Good, so the margin of error is +-5.1%

OpenStudy (anonymous):

Your a life saver. Can you help me with one more thing?

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

In a survey of 802 people, 16% said they use the internet or email more than 10 hours per week. 1. what is the margin of error for the survey? 2. Give an interval that is likely to contain the exact percent of all people who use the internet or email more than 10 hours a week.

jimthompson5910 (jim_thompson5910):

# 1 Margin of error ME = +-1/sqrt(N) x 100 ME = +-1/sqrt(802) x 100 ME = +-1/28.3196 x 100 ME = +-0.035 x 100 ME = +-3.5% So the margin of error is +-3.5% ---------------------------------------------------------- # 2 The margin of error is 3.5% So this means that our estimate 16% may be 3.5% too high or 3.5% too low. If it's 3.5% too high, then the lower bound of the interval is 16 - 3.5 = 12.5% If it's 3.5% too low, then the upper bound of the interval is 16+3.5 = 19.5% So the interval likely to contain the exact percent is (12.5%, 19.5%) You're most likely going to express the percentages in decimal form, so the interval is (0.125, 0.195)

OpenStudy (anonymous):

Thank you so much. (: This is pretty easy now. Can i ask one more favor?

OpenStudy (anonymous):

can you help me with the opposite of what we did earlier?

jimthompson5910 (jim_thompson5910):

which one do you want me to do as an example?

OpenStudy (anonymous):

number 54 & then 59 since there different, please

jimthompson5910 (jim_thompson5910):

I'm going to express everything in decimal form (so something like 5.7% is 0.057). Also, I'm going to ignore the +- for now. # 54 ME = 1/sqrt(N) 0.07 = 1/sqrt(N) 0.07*sqrt(N) = 1 sqrt(N) = 1/0.07 sqrt(N) = 14.2857 N = (14.2857)^2 N = 204.081 N = 204 So a sample size of 204 gives you a margin of error of +-7% ============================================= # 59 ME = 1/sqrt(N) 0.022 = 1/sqrt(N) 0.022*sqrt(N) = 1 sqrt(N) = 1/0.022 sqrt(N) = 45.4545 N = (45.4545)^2 N = 2066.11157 N = 2066 So a sample size of 2066 gives you a margin of error of +-2.2%

OpenStudy (anonymous):

For number 55 somehow i got the answer 16.665?

OpenStudy (anonymous):

i did it again and got 277

jimthompson5910 (jim_thompson5910):

good, your second answer of 277 is correct

OpenStudy (anonymous):

its not any of my answers tho.

OpenStudy (anonymous):

i have 278 as an answer choice

jimthompson5910 (jim_thompson5910):

that's close enough, 278 gives a margin of error of 5.99% which is very close to 6%

OpenStudy (anonymous):

Okay, Thank you so much. (:

jimthompson5910 (jim_thompson5910):

you're welcome